In a certain month, Linda earned twice as much as Jill earned. Susan earned three times as much as Jill. If Susan earned $240, how much did Linda earn?
(A) $80 (B) $160 (C) $240 (D) $360 (E) $480

Answer by palanisamy(496)

Let the earnings of Jill be x.
Given, Linda earned twice as much as Jill.
Therefore the earnings of Linda = 2x
And, Susan earned three times as much as Jill.
Therefore the earnings of Susan = 3x
Given, the earnings of Susan = $ 240
3x = 240
x = 240/3
x = 80
Therefore the earnings of Linda = 2x = 2*80 = $ 160